OpenStudy (anonymous):

How do you solve the following rational equation? Show all work. 1/x+3 +2/x+3=2/3

OpenStudy (mertsj):

\[\frac{1}{x+3}+\frac{2}{x+3}=\frac{2}{3}\] \[\frac{3}{x+3}=\frac{2}{3}\] Cross multiply and solve.

OpenStudy (anonymous):

Thanks but what would be the answer after that?

OpenStudy (mertsj):

What did you get when you cross-multiplied?

OpenStudy (anonymous):

2x-3=9

OpenStudy (mertsj):